# ![[Uncaptioned image]](extracted/5527354/head_foot/header_bar.png) CryinGAN: Design and evaluation of point-cloud-based generative adversarial networks using disordered materials – application to LiScCl-LiCoO battery interfaces | Adrian Xiao Bin Yong, and Elif Ertekin |

### v1

@@ 2.4 Evaluation of CryinGAN-generated structures
Since the best GAN architecture was found to be CryinGAN using both the fractional coordinate and bond distance discriminators, we trained a CryinGAN model () to use for further evaluation of the generated structures. The model was trained for a higher number of epochs with a shorter interval between generator trainings (see Methods section for more details). Fig. 7 shows the interface energy distribution of (relaxed) structures generated using the CryinGAN model, compared to randomly generated structures (also relaxed). The shift of the interface energy distribution to the left shows that CryinGAN has learnt to generate low-interface-energy structures. With random generation, only around 48% of the structures had low interface energy, whereas with CryinGAN, the percentage increased to around 85%. However, although CryinGAN is capable of generating low-interface-energy structures, it does not reproduce the interface energy distribution of the training dataset (yellow line). Presence of noise in the model, or mode collapse 55 where the GAN becomes biased towards generating certain types of structures only, is often the source of such differences between the generated and training distribution. We will further discuss the differences between the generated and the training structures in subsequent structural analysis.
@@
We filtered CryinGAN structures with low interface energy ( 0 J/m) for structural comparison with the training structures. We studied the structural similarity between the CryinGAN and training structures by analyzing the coordination environment (motifs) of atoms in the disordered interface region. The CrystalNNFingerprint 56, which contains the coordination likelihoods and local structure order parameters of a given atom, was calculated and averaged across structures in each dataset. We calculated the fingerprints of the cations (Li and Sc) in the interface region coordinated to the anions (Cl and O). For Li species, Table 1 shows the Euclidean distance and cosine similarity between the (average) interface fingerprint of the training and CryinGAN datasets. For comparison, we also performed the same analysis for a dataset of high-interface-energy structures (interface energy 0 J/m) that were randomly generated and relaxed. Compared to the high energy dataset, the CryinGAN dataset has a lower Euclidean distance and a higher cosine similarity to the low energy dataset, indicating that the interface Li atoms in the CryinGAN structures are more similar to the low-interface-energy training structures, than the high-interface-energy structures.
@@
0.08987
0.9942
To study the structural features in more detail, Fig. 8 shows the distribution of the most likely coordination motif of each interfacial Li atom. Although Li atoms are octahedrally coordinated in the bulk of LCO and LSC, it is found in a wide range of coordination motifs in the disordered interface region. Despite the complex structure of the interface, the Li coordination motif distribution of the CryinGAN-generated structures closely resembles that of the training structures. On the other hand, the distribution of the high-interface-energy structures is shifted towards motifs with lower coordination number (shifted upwards), indicating that the fewer bonds lead to higher energies. These results reflect the higher fingerprint similarity of the CryinGAN dataset to the training dataset as seen in Table 1, and show that CryinGAN is able to reproduce the Li coordination environment of the training structures.
Table 2 shows the Euclidean distance and cosine similarity between the interface Sc fingerprint of the CryinGAN/high energy dataset and the training dataset. Unlike the Li fingerprints, both metrics are similar for the CryinGAN and high energy datasets. This finding indicates that the average Sc coordination environment in the CryinGAN structures is not more similar to the training structures than the high energy structures. However, the Sc environments of the CryinGAN structures and the high-energy structures still exhibit differences with each other, as discussed directly below.
| Dataset |
Euclidean distance
Cosine similarity
| CryinGAN |
0.09407
0.9895
| High energy |
0.10508
0.9899
@@
Examination of the Sc coordination motif distribution shows that although the CryinGAN dataset and the high energy dataset both vary from the training dataset, they vary in different ways. Fig. 9 shows the distribution of selected coordination motifs around Sc species, focusing on the motifs that show the largest differences across the datasets. Each motif is further subdivided based on the number of O atoms in the coordination shell of Sc (the distribution of all coordination motifs is provided in ESI Fig. S10b for reference). Compared to the training dataset, the high energy dataset shows significantly higher percentages of motifs with no O atoms present, and lower percentages of motifs with O atoms. This finding indicates that the lower frequency of bonding between O (in LCO) and Sc leads to weaker interface binding and higher interface energy. In contrast, the CryinGAN dataset shows a higher frequency of Sc-O bonding than the training dataset, with higher percentages of motifs with 2 O atoms (trigonal bypyramidal and pentagonal pyramidal), and lower percentages of motifs with no O atoms (trigonal bypyramidal and octahedral). The average O bond count per Sc atom provided in ESI Fig. S11b also shows that the high energy structures have fewer Sc-O bonds, whereas the CryinGAN structures have higher number of Sc-O bonds. This analysis shows that the Sc coordination environments of the CryinGAN structures differ from the training structures due to a higher frequency of Sc-O bonding. As Sc-O bonds appear to induce low-interface-energy structures, the overpopulation of Sc-O bonds in the CryinGAN configurations is likely responsible for the lower interface energy distribution of the CryinGAN structures compared to the training dataset in Fig. 7.
We also analyzed the radial distribution functions (RDFs) of the interface Li and Sc atoms of the datasets. The RDFs confirm the trends revealed by the coordination environment analysis, in which the CryinGAN dataset shows higher similarity to the training dataset than the high energy dataset. ESI Fig. S12a-c show respectively the Sc-O, Sc-Sc, and Li-Li RDFs. For the Sc-O RDF (ESI Fig. S12a), the high energy dataset exhibits peaks with lower magnitude, reflecting its reduced degree of Sc-O bonding. On the other hand, the Sc-Sc RDF (ESI Fig. S12b) shows a higher peak around 4 Å, indicating that the high energy structures have a higher proportion of Sc cations closer together. The associated Sc-Sc repulsion raises the energy of the structures. In the Li-Li RDF (ESI Fig. S12c), a new peak appears around 2.5 Å. This short Li-Li interatomic distance again raises the energy of the structures due to Li-Li repulsion. In contrast, the CryinGAN dataset does not show any of these high-energy features, and its RDFs are similar to the training dataset.
@@ 3 Conclusions
Overall, we have shown that CryinGAN has learnt to generate disordered interface structures that not only have low interface energy, but also have good structural similarity with the training structures. This performance was achieved by training and evaluating on a fixed set of disordered materials, rather than training across compositions and space groups to generate new materials. By generating structures that can directly be compared with the training structures, it was possible to compare between model architectures and inform design choices. Including a second bond distance discriminator provided explicit information about the bond distances and improved model performance. The non-intuitive choice of using two discriminators led to better results than a single combined discriminator, despite the need for an extra tuning step. GANs also benefit from discriminators with high sampling sensitivity for generating atomic structures, but overly sensitive discriminators lead to training instability. The CryinGAN-generated interfaces exhibited the desired low-energy features of the training structures, even if the interface energy and coordination motif distributions showed differences with the training structures. Model evaluation through a fixed set of materials helps make the model development more robust and insightful, and disordered materials are sufficiently complex to be good datasets to help with this evaluation. However, generative models that incorporate symmetry constraints 57, 58 may not be able to take advantage of disordered materials for evaluation, unless the constraints can be made optional.
Amongst the options considered, the point cloud representation paired with the 1D convolutional architecture, despite its simplicity, was able to generate large, complex interface structure. Considering that the model lacks the symmetry invariances for translation, rotation, and cell repetition, its good performance is notable. However, there are some limitations to its performance. For example, we further tested CryinGAN on a dataset of amorphous silicon configurations 59 to see if it is able to generate these highly disordered structures, but found that CryinGAN unfortunately failed to generate amorphous silicon. The generated structures did not exhibit the tetrahedral coordination of amorphous silicon, and resembled randomly generated atoms (see ESI Fig. S13). It appears that to use the point cloud representation for such challenging structures, more sophisticated architectures beyond 1D convolutions are required. Graph convolution is an effective method to learn atomic structures, but implementing them in GANs requires designing a novel architecture specifically for graphs, which have already found success in VAEs and diffusion models 26, 60, 61. Autoencoders can also be combined with GANs 12, 23, where the encoded latent vectors are used as input to a GAN. Graphs are one of many options for invariant representations, among others such as smooth overlap of atomic position (SOAP) vectors 62 and the atom-centered symmetry functions (ACSFs) 63. These invariant representations should be further explored, even if reconstruction back to atomic coordinates may be needed. Reconstruction is not necessarily a barrier to their implementation, as demonstrated by Fung et al. 64 who showed that reconstruction can be achieved through gradient-based optimization using automatic differentiation. Building the next generation of generative models will rely on not only developing better architectures and representations, but also adopting better evaluation methods that provide useful feedback about model performance. We hope that the findings of this work will help advance future generative models for both ordered and disordered materials.
@@ 4 Methods
4.1 GAN and Dataset

### v2

@@ 2.5 Detailed evaluation of CryinGAN-generated interfaces
Here, we further evaluate CryinGAN to demonstrate that the generated structures are energetically and structurally similar to the training structures. We trained a CryinGAN model () for a higher number of epochs with a shorter interval between generator trainings (see Methods section for more details). Fig. 8 shows the interface energy distribution of the relaxed CryinGAN structures, compared to randomly generated structures (also relaxed). The shift of the interface energy distribution to the left shows that CryinGAN has learnt to generate low-interface-energy structures. With random generation, only around 48 % of the structures had low interface energy, whereas with CryinGAN, the percentage is around 85 %.
@@
We filtered CryinGAN structures with low interface energy ( 0 J/m2) for structural comparison with the training structures. We also compared the CryinGAN structures to a dataset of high-interface-energy structures ( 0 J/m2) that were randomly generated and relaxed. We analyzed the coordination motif fingerprints of the cations (Li and Sc) in the interface region coordinated to the anions (Cl and O), to determine the structural similarities/differences. The comparison for Li motifs showed smaller differences (see ESI Supplementary Note 3 for more details), so we focus the discussion here on Sc motifs. Table 4 shows the Euclidean distance and cosine similarity between the average interface Sc fingerprint of the CryinGAN/high energy dataset and the training dataset. The CryinGAN dataset has a lower Euclidean distance and higher cosine similarity than the high energy dataset. These results indicate that the interface Sc atoms in the CryinGAN structures are more similar to the low-interface-energy training structures, than the high-interface-energy structures.
@@
0.7074
(0.6734 to 0.7429) |
0.9316
(0.9250 to 0.9380) |
@@
Fig. 9 shows the distribution of selected Sc coordination motifs, focusing on the motifs that show the largest differences across the datasets. Each motif is further subdivided based on the number of O atoms in the coordination shell of Sc (the distribution of all coordination motifs is provided in ESI Fig. S18b for reference). Compared to the training dataset, the high energy dataset shows significantly higher percentages of motifs with no O atoms present, and lower percentages of motifs with O atoms. This finding indicates that the lower frequency of bonding between O (in LCO) and Sc leads to weaker interface binding and higher interface energy. In contrast, the CryinGAN dataset shows a higher frequency of Sc-O bonding than the training dataset, with higher percentages of motifs with 2 O atoms (trigonal bypyramidal and pentagonal pyramidal), and lower percentages of motifs with no O atoms (trigonal bypyramidal and octahedral). The average O bond count per Sc atom provided in ESI Fig. S19b also shows that the high energy structures have fewer Sc-O bonds, whereas the CryinGAN structures have higher number of Sc-O bonds. This analysis shows that although the Sc coordination environments of the CryinGAN structures differ from the training structures due to a higher frequency of Sc-O bonding, the Sc-O bonds are still a feature of the low-interface-energy structures. We also analyzed the RDFs of the interface Li and Sc atoms. The RDFs confirm the trends revealed by the coordination motif analysis, in which the CryinGAN dataset shows higher similarity to the training dataset than the high energy dataset (see ESI Supplementary Note 4 for more details).
2.6 Overall generative model comparisons and insights
The Dismai-Bench metrics of CryinGAN across all datasets are listed in Section 2.3. For the disordered interfaces, CryinGAN outperforms both U-Net diffusion models, despite that all three models use coordinate-based representations, and diffusion models are often reported to outperform GANs in image synthesis 60, 61, 28. CryinGAN performs competitively with the graph diffusion models for the disordered interfaces, which is unexpected considering the lack of invariances and graph convolutions in CryinGAN. Nonetheless, the limited expressive power of point clouds does introduce challenges when CryinGAN is tasked with generating amorphous Si structures. However, CryinGAN is still able to generate the disordered alloy structures and reproduce the SRO distributions, while the other coordinate-based diffusion models struggled with this task. Apart from the 300 K dataset with narrow SRO, CryinGAN demonstrated similar metrics to the graph diffusion models (refer to Table 3). The results of this simple point-cloud-based GAN highlights the benefits and importance of robust generative model evaluation.
An overall ranking of all generative models based on their general performance on each Dismai-Bench dataset is visualized using a spider chart as shown in Fig. 10. The U-Net diffusion models struggled with most tasks in Dismai-Bench, while performing the best on the disordered interface dataset. Between the two, CrysTens performed marginally better on the disordered interfaces, probably as a result of the inclusion of distance matrices (albeit at a memory cost that scales with O(), being the number of atoms). Also, both UniMat and CrysTens require data augmentation to compensate for their lack of all invariances, further increasing their memory requirements by orders of magnitude. We expect that the reason for the better performance on the disordered interface dataset is that the interface structures have sorted atom orderings that make learning easier for these models with image-/video-like representation. For example, the first 72 atoms are always Li, with the first 9 Li atoms always in the interface region, the next 9 Li atoms in the top layer of LCO, etc. In contrast, the other datasets do not have such well-sorted atom orderings. When CrysTens and UniMat were trained on disordered interface structures with randomized atom orderings, they were unable to generate the interface structures (see ESI Fig. S20). On the other hand, CryinGAN is permutationally invariant, so it does not rely on the sorted atom orderings to perform well on the interface dataset, and was able to generate the crystalline alloy structures.
The graph diffusion models demonstrated the best performance overall. Comparing the two models, CDVAE performed better on the amorphous Si and the 300 K alloy datasets, whereas DiffCSP performed better on the 1500 K alloy and interface datasets. CDVAE stands out as the only model that is able to generate structures that are nearly fully relaxed. In use cases where relaxation/post-processing is not feasible/practical (e.g., high-throughput materials discovery), CDVAE may show some advantages here. Although the graph models were successful in performing diffusion on atomic coordinates alone, they suffered when performing joint diffusion with atomic species or lattice parameters. When using atomic species diffusion, CDVAE was not able to generate the compositions correctly even though all Dismai-Bench datasets have fixed compositions. When using lattice diffusion, DiffCSP generated structures of lower quality. Such issues may be less severe when training on datasets with smaller number of atoms or lattice lengths, but future models can still improve upon their joint diffusion performance. To this end, Dismai-Bench provides a means for testing and evaluating different features/architectures (e.g., training DiffCSP with teacher forcing of the lattice parameters; see ESI Fig. S6).
Despite the superior performance of graph models on Dismai-Bench, there are also reports that less expressive generative models (e.g., UniMat 29, language models 31, 32, 33) can show comparable or even better performance in discovering novel materials than graph models. These seemingly conflicting observations suggest a need to reconsider material discovery strategies. Materials discovery can be formulated as an exploration-exploitation problem. Graph models are superior at learning structural patterns, and should likely learn and exploit more of the training distribution during generation. On the other hand, less expressive models learn more noisy distributions, and can readily explore outside of the training distribution during generation. Considering the relatively small size of training datasets ( for the MP-20 dataset 26) vs. the material space of possible compounds ( possible combinations of quaternary compounds 62), it is arguable that exploration is more important that exploitation in materials discovery. However, good performance for materials discovery should not be confused with the ability to learn from training structures. While less expressive models may do well for the discovery of small structures, they may struggle when the size/complexity of the structures increases, as suggested here.
As the number of atoms and/or structural complexity of training samples increases, the importance of symmetry invariances and the expressive capability of generative models also grows. Graphs serve as one type of invariant representation among alternatives like smooth overlap of atomic positions (SOAP) vectors 63 and atom-centered symmetry functions (ACSFs) 64. Exploring these invariant representations further is crucial, even if reconstruction back to atomic coordinates is necessary. Importantly, reconstruction does not necessarily pose a fundamental barrier, as demonstrated by Fung et al. 65, who achieved it through gradient-based optimization using automatic differentiation. The Dismai-Bench findings underscore the need for higher expressive power to model more complex structures, despite potential increases in computational and memory requirements. As generative modeling advances towards larger datasets and more complex structures, the development of models capable of parallelization will be a critical research direction.
In this work, we introduced Dismai-Bench as a novel method for assessing generative models and obtaining valuable insights into their performance. The development of generative models, and machine learning models in general, is iterative and often counterintuitive, as evidenced by the development of CryinGAN. Surprising findings include instances where an older generation GAN architecture has outperformed newer diffusion architectures, and that using two separate discriminators has yielded better results than using a single discriminator. Dismai-Bench proves effective in evaluating a generative model’s capability to learn intricate structural patterns, particularly those present in disordered materials. Consequently, Dismai-Bench may contribute to improving generative models not only for ordered materials but also for disordered ones. However, Dismai-Bench’s scope is limited in that it primarily assesses the accuracy of atomic positions, and does not evaluate a model’s proficiency in learning compositions or lattice parameters, because Dismai-Bench operates by fixing the composition and lattice while varying atomic positions. Moreover, generative models incorporating symmetry constraints 66, 67 may not leverage Dismai-Bench for evaluation unless these constraints can be disabled. Dismai-Bench marks a significant initial stride towards robustly evaluating generative models, and we anticipate that new complementary benchmarks and datasets will emerge in the future.
@@ 3 Conclusions
We developed a new benchmark for generative models, Dismai-Bench, which evaluates models on datasets of large, disordered materials exhibiting different degrees of structural and configurational disorder. Instead of training across different compositions and space groups, models are trained on datasets with fixed composition and structure type. By fixing the material system, Dismai-Bench circumvents the challenges of evaluating models based on newly generated materials that cannot be verified. Graph diffusion models (CDVAE & DiffCSP) were found to outperform coordinate-based diffusion models (CrysTens & UniMat) on Dismai-Bench, due to the invariant nature and higher expressive power of graphs. Additionally, we introduced CryinGAN, a novel GAN based on point clouds, that was developed by evaluating candidate architectures through direct comparisons between training and generated structures. Despite its simple architecture without symmetry invariances or complex components, CryinGAN outperformed the coordinate-based diffusion models and demonstrated competitiveness with the graph diffusion models. Dismai-Bench provides meaningful evaluation for comparing between architectures, understanding model strengths and weaknesses, and ultimately informing design choices. Building the next generation of generative models will rely on not only developing better architectures and representations, but also adopting better evaluation methods. We hope that this work will help advance future generative models for both ordered and disordered materials, and inspire the development of other new innovative benchmarks.
@@ 4 Methods
4.1 Dismai-Bench datasets
Dismai-Bench uses a total of six datasets. Each dataset contains a total of 1,500 structures, split into 80 % training and 20 % validation data.
4.1.1 Fe60Ni20Cr20 austenitic stainless steel
A cluster expansion Monte Carlo approach 35, 36 was used to generate the datasets of FCC Fe60Ni20Cr20 austenitic stainless steels. The CE model operates as a generalized Ising model 68 to describe the formation energy as a function of configuration. We adapted the CE model from ref. 36. A CE model with seven chemical dimers and three spin dimers was used to fit a dataset of FCC Fe-Ni-Cr alloys that were generated from spin-polarized DFT calculations. The CE model was fit using the least absolute shrinkage and selection operator (LASSO) and 10-fold cross-validation (CV) to determine the effective cluster interactions (ECI) values of the clusters. A comparison between the CE-calculated and DFT-calculated formation energies is shown in ESI Fig. S21a, and the ECI values obtained by the LASSO CV fit are shown in ESI Fig. S21b.
To generate the Dismai-Bench alloy structures, only the chemical terms of the CE model were used to obtain structures in a non-magnetic state. A 444 conventional FCC supercell, containing 256 atoms (154 Fe atoms, 51 Ni atoms, 51 Cr atoms), was used. We performed MC simulations in the canonical ensemble based on the Metropolis-Hastings algorithm 69. Kawasaki dynamics 70 for atom swaps was applied to ensure that the composition of the system remained fixed. The alloy structures were sampled at temperatures of 300 K and 1,500 K, where 10 independent MC simulations were initialized for each temperature. For every MC simulation, 50,000 passes were performed, and a MC snapshot was saved every 10 passes. For each temperature, 1,500 structures were assembled (without any restriction on SRO distribution) as the wide SRO dataset. Another 1,500 structures with SRO distribution within 0.1 of the average SRO values were filtered as the narrow SRO dataset. Comparisons of the SRO between the Fe60Ni20Cr20 alloy modelled in Dismai-Bench and the original Fe56Cr21Ni23 alloy modelled in ref. 36 is shown in ESI Fig. S21c. The Dismai-Bench alloy shows qualitatively similar SRO trends to the original alloy that was modelled with magnetism and a larger set of clusters.
4.1.2 Amorphous silicon
We adapted the amorphous silicon dataset from ref. 38. The original data consists of a 100,000-atom amorphous silicon structure generated through melt-quench molecular dynamics simulation 38. We sliced the structure into smaller blocks with lattice parameters corresponding to 256-atom amorphous silicon structures. The lattice lengths of the blocks were calculated by linearly scaling the lattice lengths of the 100,000-atom structure to 256 atoms. The blocks were sliced at different locations to obtain a total of 1,500 blocks. Blocks with < 256 atoms had atoms added at random to low density regions, and blocks with > 256 atoms had atoms removed at random from high density regions, until all blocks had 256 atoms. Atoms were added/removed to the boundary of the blocks only (where they were sliced). The density was calculated by dividing each face of a block into 22 regions and counting the number of atoms in each region. The 1,500 blocks were relaxed using a pre-trained SOAP-GAP 39 machine learning interatomic potential for Si. The structures were optimized using a conjugate gradient algorithm 71 through the atomic simulation environment (ASE) package 72. Only the atomic positions were allowed to relax, and the relaxations were stopped when the force on each atom was below 0.05 eV/Å.
4.1.3 Li3ScCl6(100)-LiCoO2(110) battery interface
To construct the interface structures, we chose the orientations of LCO(110), a Li fast-diffusing plane 73, and LSC(100), a representative plane. The surfaces of the LSC(100) slab are polar, so half of the Cl atoms were moved from one surface to the other to neutralize the polarity (resulting in ‘Tasker Type 2b’ surfaces 74, 75). Lattice matching between the two slabs was carried out using the MPInterfaces package 76, which implements the lattice matching algorithm proposed by Zur et al. 77. The configuration of the lattice-matched interface is given in ESI Table S4, where the average lattice mismatch is 2.17 %.
In a preliminary test calculation, the interface was constructed by simply placing the LCO(110) and LSC(100) slabs in contact with each other, and we obtained a DFT-relaxed structure similar to that depicted in Fig. 1. An interface region with disordered LSC atoms formed with a thickness of around 5 Å. Using this structure as reference, we then generated random interface structures using the CALYPSO package 78, 79, 80. We used LCO(110) and LSC(100) slabs with 4 and 7 layers respectively. For each structure, an interface region thickness was randomly chosen between 4 and 6 Å, then the region was randomly populated with 3 formula units of LSC (9 Li, 3 Sc, and 18 Cl atoms). A random lateral displacement (in-plane direction parallel to the interface) was also applied to the LSC slab. A vacuum spacing of 14 Å was included in all the interface structures. Each interface structure has a total of 264 atoms.
We relaxed the randomly generated interface structures with the M3GNet interatomic potential. Only the atom positions were allowed to relax, where the lateral lattice vectors were fixed to the optimized values of the LCO slab (the elastic moduli of LCO 81 is significantly larger than LSC 82). Similar to the amorphous Si dataset, the structures were optimized using a conjugate gradient algorithm through the ASE package, and the relaxations were stopped when the force on each atom was below 0.05 eV/Å. The normalized interface energies, , of the relaxed structures were calculated using the following equation:
where is the total energy of the interface in eV, is the total number of atoms, and is the interface area. The interface energies are normalized such that structures with interface energy 0 J/m2 are considered to be low-interface-energy structures. 1,500 low-interface-energy structures were filtered from the relaxed structures to form the dataset.
4.2 Generative models
4.2.1 CDVAE
We modified CDVAE such that atomic species denoising becomes an optional feature, since atomic species denoising caused CDVAE to generate incorrect compositions when trained on Dismai-Bench datasets. All CDVAE models were trained without atomic species denoising. The same hyperparameters, optimizer, and learning rate scheduler as those used for the MP-20 dataset 26 were applied. All models were trained using a batch size of 8 for 1,000 epochs. Disordered interface and alloy structures were generated by running Langevin dynamics using 100 steps per noise level. For the amorphous Si structures, we were able to match the energy distribution of generated structures to the training structures (refer to example in ESI Fig. S22), so we used 6-7 steps per noise level based on whichever setting gave the best match.
4.2.2 DiffCSP
We used 4 layers and 256 hidden states for all DiffCSP models. The exponential noise scheduler parameter, 27, was set to 0.05 for amorphous Si, and 0.1 for disordered interfaces and alloys. The weight of the lattice cost was set to 0 for all models (i.e., no lattice diffusion). All other hyperparameters were set to the default values, and we used the same optimizer and learning rate scheduler as the original implementation 27. All models were trained using a batch size of 4 for 1,000 epochs. All structures were generated using step size and 1000 time steps. Although we added the option to train DiffCSP with teacher forcing of the lattice parameters, this feature was not used for the DiffCSP models benchmarked in Section 2.3, which did not use any lattice diffusion.
4.2.3 CrysTens
CrysTens uses the Imagen 83 model, a cascaded diffusion model consisting of a base U-Net that generates a lower resolution image, followed by a super-resolution U-Net that upsamples the lower resolution image to a higher resolution. We used 64 base channels for both U-Nets, where the first U-Net generates 6464 images and the second U-Net upsamples them to the full-size CrysTens images. The optimizer and all other hyperparameters were set to be the same as the original CrysTens implementation 28. Each U-Net was trained using a batch size of 8 for 150,000 training steps. Structures were generated using 100 time steps, where the energy distribution of the generated structures had already converged at this setting (refer to ESI Fig. S23). The CrysTens images were reconstructed into atomic structures using the ground truth lattice parameters. Althought the generated lattice lengths and angles from the CrysTens images were not used, their MAEs were small, around 0.02 Å and 0.02∘ respectively. The composition accuracy of the generated structures was 100 %. The atomic coordinates of each atom was determined by constructing directional graphs using the coordinate pixels and the pairwise , , and pixels, as described in ref. 28. Then, the coordinate predictions from the directional graphs were averaged. However, unlike the original implementation, we did not perform -means clustering on the averaged atomic coordinates (and atomic species), since this was a post-processing step intended to manually "denoise" the reconstructed structures using an arbitrary choice of .
For CrysTens, we augmented the disordered interface dataset by applying random permutations (shuffling) to the atoms in the interface region. CrysTens was found to perform poorly when data augmentation was performed by permutating all atoms (see ESI Fig. S20). Each structure in the interface dataset was constructed with the same LSC and LCO slabs, as well as three formula units of LSC randomly generated in the interface region. As a result, the ordering of the atoms in the slabs is consistent across all structures, making it easier for CrysTens to learn these structures. Therefore, we only shuffled the ordering of the atoms in the interface region, where each structure was augmented 49 times. For amorphous Si and the alloys, which have no consistent atom orderings in their datasets, CrysTens was unable to generate meaningful structures even with data augmentation (see Fig. 3 and Fig. 4).
4.2.4 UniMat
The original UniMat 29 code is unfortunately not openly available, so we used an open-access implementation 53 of the 3D U-Net model 54 that the UniMat model was repurposed from. We used the hyperparameters as listed in ESI Table S5 for the U-Net. 44 video frames were used for the UniMat representation, which is the smallest frame size compatible with U-Net model configuration. The Adam optimizer 84 was used with learning rate = , , and . The U-Net was trained using a batch size of 8 for 150,000 training steps. Structures were generated using 100 time steps, where the energy distribution of the generated structures had already converged at this setting (refer to ESI Fig. S24). We did not include the lattice parameters in the UniMat representation, and simply used the ground truth lattice parameters to reconstruct the structures (i.e., no lattice diffusion). The average composition accuracy of the generated structures was 99.5 %. Structures with incorrect compositions were considered failed structures and rejected. Similar to CrysTens, we augmented the disordered interface dataset by shuffling the atoms in the interface region, augmenting 49 times per structure. For amorphous Si and the alloys, UniMat was unable to generate meaningful structures even with data augmentation (see Fig. 3 and Fig. 4).
4.2.5 CryinGAN